OpenStudy (anonymous):

(a) find the Maclaurin series for the following function (1). (b) use the series in (a) to evaluate (2)

OpenStudy (anonymous):

(1)\[\left[\begin{matrix}1 & x=0 \\ (e ^{x}-1)/x& x \neq0\end{matrix}\right]\] (2)\[\sum_{n=1}^{\infty}n/(n+1)!\]